Determine maximum diameter of hole that can be punched through a plate of 10mm thick,if the punch can withstand a compressive stress of 500mpa and ultimate shear stress of 200mpa


1
Expert's answer
2021-11-17T08:20:03-0500


Given data:

"t=10mm"

"\sigma_c=500 Mpa"

"\tau=200 Mpa"

Solution:

as we know that

"\sigma_c\times A_c = \tau\times A_s"

"\sigma_c\times\pi/4 D^2=\tau\times\pi D t"

"500/4 (D)= 200\times 10"

"D= 16 mm"

This is the required diameter of the hole